OpenStudy (anonymous):

Admission to the movies costs $8 for adults and $4 for kids. The total bill for your entire extended family was $72. This scenario can be modeled by the equation 8a + 4k = 72. Which is not a possible combination of adults and kids? Answer 3 adults and 12 kids 7 adults and 4 kids 8 adults and 2 kids 5 adults and 6 kids

Parth (parthkohli):

The last one, as: \(\Large \color{Midnightblue}{\rightarrow 8(5) + 4(6) = 40 + 24 = 64 \ne 72 }\)

OpenStudy (anonymous):

wat